tblanks94:

Using your calculator, find the money accumulated if I invested $65,000 at 7.3% interest for 10 years compounded continuously.

jhonyy9:

use the rule of 3 simple and write 65000 mean 100% x mean 7,3% x = (7,3 *65000 ) / 100 = ? x will be the interest for one year

Mercury:

continuous compounding does not use the same formula as regular compound interest A = Pe^(rt) where A is the final amount, P is the principal (65,000), r is the interest rate (0.073) and t is time (10 years). evaluate for A.